if ratio of age of A:B is 6:5and the age of each C and D is 9/10 times that of B.age of F is less than A but greater tha B.the ratio of age of b and E is 2:3 also age of A is 3years less than E.what is the ratio of age of A and F,if all ages are in integer?
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
if ratio of age of A:B is 6:5
let x = the first multiplier
then
6x = A's age
5x = B's age
and the age of each C and D is 9/10 times that of B.
C =.9B
D = .9B
age of F is less than A but greater than B.
F < A
F > B
the ratio of age of B and E is 2:3
let y = the 2nd multiplier
2y = B's age
3y = E's age
also age of A is 3years less than E.
A = E-3
what is the ratio of age of A and F, if all ages are in integer?
:
We can reason this out, assume the first multiplier is 2, then
6(2) = 12 yrs is A's age
5(2) = 10 yrs is B's age
F < A & F > B therefore
F = 11 yrs old
The A:F ratio then 12:11
:
:
We don't need the other ages but we can find them easily
C & D = .9(10) = 9 yrs old
2y = 10 (B's age)
y = 5 is the 2nd multiplier
E = 3(5) = 15 yrs is E's age
which agrees with the statement
"age of A is 3years less than E."
A = 12 yrs, E = 15 yrs
:
Summarize: A = 12; B = 10; C = 9; D = 9; E = 15; F = 11